# On the Fragility of Active Learners

### 23 Mar 2024 v1

@@ 4 Comparison Methodology
We run an experiment till the size of the labeled set has grown to $5000$ instances. This which implies $T=(5000-200)/200=24$ iterations for the batch/seed size setting of $(200,200)$, and similarly $9$ iterations for the $(500,500)$ setting.
@@ 6 Results
  • $\mathcal{D}\in\{imdb,agnews,dbpedia-5,pubmed,sst-2\}$ is a query strategy.
@@ 6.4 The best combinations
Table 1 shows the top-$10$ combinations of pipelines and query strategies for different regimes. Not surprisingly, wrt pipelines, RoBERTa shows the strongest performance. Somewhat surprisingly, the next best pipeline is LinearSVC-MPNet (shortened to LinSVC_MPNet) and not a RF-based pipeline. Our guess as to why this is the case is that RF’s hyperparamṡearch space (detailed in Section 0.A.4.2) was not comprehensive enough, while MPNet transformed the data to be nearly linearly separable, which enabled the high-bias LinearSVC to quickly find a good optima. Sadly, no single query strategy dominates the QS column, and thus we are offered no insight as to what strategy pairs best with a pipeline. Somewhat noteworthy is the fact that the random strategy is quite competitive for the LinearSVC-MPNet pipeline, but not for RoBERTa.
@@
(0-1250] 2. RoBERTa real 873.64 7. LinSVC_MPNet real 794.69
(0-1250] 3. RoBERTa margin 873.36 8. LinSVC_MPNet margin 793.08
(0-1250] 4. RoBERTa cal 871.92 9. LinSVC_MPNet cal 790.42
(0-1250] 5. RoBERTa random 862.99 10. LinSVC_MPNet dal 789.60
@@
(1250-2500] 2. RoBERTa dal 984.40 7. LinSVC_MPNet real 907.98
(1250-2500] 3. RoBERTa real 981.07 8. LinSVC_MPNet margin 907.39
(1250-2500] 4. RoBERTa margin 976.07 9. LinSVC_MPNet dal 900.13
(1250-2500] 5. RoBERTa random 974.04 10. LinSVC_MPNet cal 899.12
@@
(2500-3750] 2. RoBERTa dal 900.93 7. LinSVC_MPNet margin 836.57
(2500-3750] 3. RoBERTa real 900.83 8. LinSVC_MPNet real 833.95
(2500-3750] 4. RoBERTa cal 897.16 9. LinSVC_MPNet dal 828.73
(2500-3750] 5. RoBERTa random 891.58 10. LinSVC_MPNet cal 827.24
@@
(3750-5000] 2. RoBERTa margin 1087.37 7. LinSVC_MPNet random 1010.88
(3750-5000] 3. RoBERTa dal 1085.06 8. LinSVC_MPNet real 1008.67
(3750-5000] 4. RoBERTa cal 1076.79 9. LinSVC_MPNet dal 1004.44
(3750-5000] 5. RoBERTa random 1075.20 10. LinSVC_MPNet cal 999.97
@@ 0.A.3 Reproducibility Experiments
In all cases, we have attempted to replicate the original settings, e.g. same train/development/test data split, model type, seed/batch sizes, number of AL iterations as shown in Table 3. For CAL, REAL, we report the F1-macro scores on agnews, in which classes are evenly distributed, instead of the accuracy provided in the original papers. For DAL, we use the dataset cola [^8] https://nyu-mll.github.io/CoLA/ and utilise the Hugging Face library to finetune BERT (while the original work employs TensorFlow [^9] https://www.tensorflow.org/, but we use equivalent settings). Figure 9 shows a comparison between our results and the reported ones in these papers. Despite some minor differences in the setups, we observe that these AL methods work as described in their respective papers in these settings.
@@
Figure 9: Comparison between published results and ours with the same settings for CAL, REAL, DAL.

### 11 Apr 2024 v2

@@ 4 Comparison Methodology
We run an experiment till the size of the labeled set has grown to $5000$ instances. This implies $T=(5000-200)/200=24$ iterations for the batch/seed size setting of $(200,200)$, and similarly $9$ iterations for the $(500,500)$ setting.
@@ 6 Results
  • $\mathcal{D}\in\{imdb,agnews,dbpedia-5,pubmed,sst-2\}$ is a dataset.
@@ 6.4 The best combinations
Table 1 shows the top-$10$ combinations of pipelines and query strategies for different regimes. Not surprisingly, wrt pipelines, RoBERTa shows the strongest performance. Somewhat surprisingly, the next best pipeline is LinearSVC-MPNet (shortened to LinSVC_MPNet) and not a RF-based pipeline. Our guess as to why this is the case is that RF’s hyperparam. search space (detailed in Section 0.A.4.2) was not comprehensive enough, while MPNet transformed the data to be nearly linearly separable, which enabled the high-bias LinearSVC to quickly find a good optima. Sadly, no single query strategy dominates the QS column, and thus we are offered no insight as to what strategy pairs best with a pipeline. Somewhat noteworthy is the fact that the random strategy is quite competitive for the LinearSVC-MPNet pipeline, but not for RoBERTa.
@@
2. RoBERTa real 873.64 7. LinSVC_MPNet real 794.69
3. RoBERTa margin 873.36 8. LinSVC_MPNet margin 793.08
4. RoBERTa cal 871.92 9. LinSVC_MPNet cal 790.42
5. RoBERTa random 862.99 10. LinSVC_MPNet dal 789.60
@@
2. RoBERTa dal 984.40 7. LinSVC_MPNet real 907.98
3. RoBERTa real 981.07 8. LinSVC_MPNet margin 907.39
4. RoBERTa margin 976.07 9. LinSVC_MPNet dal 900.13
5. RoBERTa random 974.04 10. LinSVC_MPNet cal 899.12
@@
2. RoBERTa dal 900.93 7. LinSVC_MPNet margin 836.57
3. RoBERTa real 900.83 8. LinSVC_MPNet real 833.95
4. RoBERTa cal 897.16 9. LinSVC_MPNet dal 828.73
5. RoBERTa random 891.58 10. LinSVC_MPNet cal 827.24
@@
2. RoBERTa margin 1087.37 7. LinSVC_MPNet random 1010.88
3. RoBERTa dal 1085.06 8. LinSVC_MPNet real 1008.67
4. RoBERTa cal 1076.79 9. LinSVC_MPNet dal 1004.44
5. RoBERTa random 1075.20 10. LinSVC_MPNet cal 999.97
@@ 0.A.3 Reproducibility Experiments
In all cases, we have attempted to replicate the original settings, e.g. same train/development/test data split, model type, seed/batch sizes, number of AL iterations as shown in Table 3. For CAL, REAL, we report the F1-macro scores on agnews, in which classes are evenly distributed, instead of the accuracy provided in the original papers. For DAL, we use the dataset cola [^8] https://nyu-mll.github.io/CoLA/ and utilise the Hugging Face library to finetune BERT (while the original work employs TensorFlow [^9] https://www.tensorflow.org/, but we use equivalent settings). Figure 9 shows a comparison between our results and the reported ones in these papers [21, 5, 11] for CAL, REAL, DAL, respectively. Despite some minor differences in the setups, we observe that these AL methods work as described in their respective papers in these settings.
@@
Figure 9: Comparison between published results in [21, 5, 11] and ours with the same settings for CAL, REAL, DAL.